OpenStudy (anonymous):

Susan is reading a book. The number of pages before the page she is reading is equal to 1/10 the number of pages after. If she reads 90 more pages, the number of pages before that page equals the number of pages after. Find the total number of pages of the book. i even don't understand what does the question is saying.... @Peter14

OpenStudy (anonymous):

let a be the number of pages she has read and b be the number of pages left

OpenStudy (anonymous):

set up a system of equations...

OpenStudy (anonymous):

can we use one unknown? or, just go on?

OpenStudy (anonymous):

90/100 = 0.9 so, 0.9*10 =9 9+1 = 10 so, 10+90 = 100 100+100 = 200 so, the book has 200 pages.

OpenStudy (anonymous):

wow, so many people...

OpenStudy (anonymous):

@ryanvarghese12790 sorry the answer is more than 200..

OpenStudy (anonymous):

I am hoping this is right.

OpenStudy (anonymous):

sorry, then I am out

OpenStudy (anonymous):

10*a=b (a+90)=(b-90)

OpenStudy (anonymous):

@ryanvarghese12790 never mind :)

OpenStudy (anonymous):

@Peter14 ok just ignore me xdd. go on then. let me see...

OpenStudy (anonymous):

Look ! We have a Proportion ! we have 1/10 means difference is 10-1=9 Then : 1___________x 10______-y 9_______90 Have : 10*10=100 and : 1*10=10 So sum them : 100+10=110 Got it ?:) @kryton1212

OpenStudy (anonymous):

@E.ali i am sorry. the answer is not 110.....

OpenStudy (anonymous):

Ohhhh .... yea .... I have a bigggggggggg fall ... Sorry !!! Just a min !...

OpenStudy (anonymous):

@E.ali never mind :D no need to apologize :)

OpenStudy (anonymous):

10*a=b (a+90)=(b-90) ==>> a+180 = b a+180 = 10a

OpenStudy (anonymous):

@kryton1212 you got it from here

OpenStudy (anonymous):

Yeah ! Look if we reduce from a thing another will be more !

OpenStudy (anonymous):

well.....

OpenStudy (anonymous):

For example : a+b=8 If have a-4 have b+4 : a-4+b+4=8 Got it ?

OpenStudy (anonymous):

a=20 b=200 and include the first page 20+200+1=221?

OpenStudy (anonymous):

ohh thank you @E.ali and @Peter14 so much

OpenStudy (anonymous):

sure, include the first page

OpenStudy (anonymous):

Yea !

OpenStudy (anonymous):

thanks ;D

OpenStudy (anonymous):

;)